In a basketball game Kyle scored 4 more than 1/5 as many points as Kurt. If Kyle scored 12 points, how many points did Kurt score?
Answer by EMStelley(208)

Call the number of points Kurt scored p. Then Kyle scored (1/5)p+4. Since Kyle scored 12, we have
(1/5)p+4=12
(1/5)p=8
p=40
So Kurt scored 40 points.